OpenStudy (alexandervonhumboldt2):

Simplify:

OpenStudy (alexandervonhumboldt2):

((a^2-b^2-c^2-2bc)*(a+b-c))/((a+b+c)*(a^2-b^2+c^2+2ac))

OpenStudy (alexandervonhumboldt2):

Hint: Use foolowing statementts. a^3+b^3+c^3-3abc=(a+b+c)^3-3*(a+b+c)(ab+bc+ca) and a^2+b^2+c^2-ab-bc-ca=(a+b+c)^2-3(a+b+c)*(ab+bc+ca).

OpenStudy (alexandervonhumboldt2):

this is not immediate so type your idea, when i'll be online i'll see what you typed.